Two objects attract each other with a gravitational force of magnitude 9.60 10-9 N when separated by 19.9 cm. If the total mass of the objects is 5.09 kg, what is the mass of each? (a) heavier mass (b) lighter mass
Answer #1

The force of attraction between the two bodies is F= \frac{Gm_{1}m_{2}}{r^{2}}

for heavier and lighter mass we write m1 and m2 = 5.09-m1

Then from deffinition 9.6x10^{-9}= \frac{Gm_{1}(5.09-m_{1})}{0.199^{2}}

  m_{1}(5.09-m_{1})= 9.6x10^{-9}x0.199^{2}/6.67x10^{-11}

5.09m_{1}-m_{1}^{2} = 5.70

  m_{1}^{2} -5.09m_{1}+5.70=0

From quadratic equation we have x=\frac{-b\pm \sqrt{b^{2}-4ac}}{2a}

  \Rightarrow m_{1}=\frac{5.09\pm \sqrt{5.09^{2}-4x1x5.7}}{2}

m1  = (5.09+1.76)/2 and m2 = (5.09-1.76)/2

m1 = 3.425kg

m2 = 1.665kg